# Hui Xiong2, Bin Yang1

### v1

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE Ocsvm IF PCA HBOS
CICIDS Aff-P 0.667 0.563 0.586 0.553 0.531 0.533 0.566 0.569 0.575 0.579 0.543 0.530 0.548 0.541 0.538
Aff-R 0.959 0.781 0.892 0.926 0.593 0.882 0.783 0.785 0.799 0.794 0.156 1.000 0.672 0.724 0.545
Aff-F 0.787 0.654 0.708 0.692 0.560 0.664 0.657 0.660 0.669 0.669 0.243 0.693 0.604 0.619 0.542
CalIt2 Aff-P 0.742 0.650 0.703 0.617 0.645 0.571 0.691 0.667 0.668 0.616 0.560 0.652 0.539 0.688 0.620
Aff-R 0.955 0.975 0.963 0.959 0.838 0.894 0.932 0.976 0.976 0.984 0.617 0.982 0.321 0.869 0.969
Aff-F 0.835 0.780 0.812 0.751 0.729 0.697 0.794 0.793 0.793 0.757 0.587 0.783 0.402 0.768 0.756
Credit Aff-P 0.618 0.611 0.559 0.513 0.521 0.488 0.610 0.612 0.605 0.608 0.560 0.556 0.658 0.564 0.533
Aff-R 0.956 0.952 0.984 0.935 0.865 0.898 0.953 0.955 0.948 0.952 0.562 0.997 0.612 0.959 0.999
Aff-F 0.750 0.744 0.713 0.663 0.650 0.632 0.744 0.746 0.738 0.742 0.561 0.714 0.634 0.710 0.695
GECCO Aff-P 0.832 0.808 0.735 0.633 0.690 0.567 0.810 0.831 0.808 0.793 0.836 0.499 0.647 0.646 0.620
Aff-R 0.998 0.998 0.979 0.786 0.903 0.872 0.997 0.995 0.997 0.992 0.810 1.000 0.315 1.000 0.827
Aff-F 0.908 0.893 0.839 0.701 0.782 0.687 0.894 0.906 0.893 0.882 0.823 0.666 0.424 0.785 0.708
Genesis Aff-P 0.835 0.728 0.822 0.683 0.749 0.659 0.780 0.763 0.772 0.724 0.759 0.512 0.673 0.691 0.564
Aff-R 0.966 0.974 0.972 0.996 1.000 0.943 0.968 0.974 0.959 0.971 0.976 1.000 0.951 0.991 1.000
Aff-F 0.896 0.833 0.891 0.810 0.856 0.776 0.864 0.856 0.856 0.829 0.854 0.677 0.788 0.814 0.721
MSL Aff-P 0.599 0.578 0.566 0.562 0.549 0.576 0.589 0.584 0.577 0.584 0.521 0.497 0.502 0.538 0.520
Aff-R 0.966 0.975 0.951 0.618 0.933 0.874 0.973 0.952 0.975 0.948 0.781 0.902 0.697 0.914 0.982
Aff-F 0.740 0.726 0.710 0.588 0.692 0.694 0.734 0.724 0.725 0.723 0.625 0.641 0.584 0.678 0.680
NYC Aff-P 1.000 0.639 0.520 0.551 0.751 0.769 0.750 0.719 0.814 0.817 0.529 0.500 0.481 0.516 0.525
Aff-R 0.989 0.965 0.999 0.988 0.988 0.980 0.843 0.843 0.843 0.821 0.989 1.000 0.989 0.997 0.946
Aff-F 0.994 0.769 0.684 0.708 0.853 0.862 0.794 0.776 0.828 0.819 0.689 0.667 0.648 0.680 0.675
PSM Aff-P 0.808 0.734 0.765 0.622 0.600 0.538 0.762 0.739 0.777 0.762 0.776 0.652 0.904 0.712 0.621
Aff-R 0.918 0.941 0.966 0.868 0.871 0.932 0.939 0.948 0.893 0.942 0.649 0.447 0.472 0.692 0.700
Aff-F 0.859 0.825 0.854 0.725 0.710 0.682 0.842 0.831 0.831 0.843 0.707 0.531 0.620 0.702 0.658
SMD Aff-P 0.773 0.755 0.736 0.527 0.607 0.510 0.745 0.748 0.761 0.762 0.889 0.649 0.801 0.680 0.557
Aff-R 0.938 0.948 0.943 0.956 0.895 0.998 0.938 0.970 0.940 0.946 0.291 0.866 0.513 0.807 0.722
Aff-F 0.847 0.840 0.827 0.679 0.724 0.675 0.831 0.845 0.841 0.844 0.439 0.742 0.626 0.738 0.629
Contextual4.9 Aff-P 0.838 0.505 0.821 0.503 0.517 0.490 0.552 0.669 0.773 0.563 0.861 0.619 0.861 0.488 0.487
Aff-R 0.812 0.814 0.761 0.776 0.792 0.779 0.837 0.883 0.806 0.908 0.674 0.769 0.582 0.463 0.547
Aff-F 0.825 0.623 0.790 0.611 0.625 0.602 0.666 0.761 0.789 0.695 0.756 0.686 0.694 0.475 0.515
Contextual7.2 Aff-P 0.907 0.508 0.826 0.519 0.498 0.496 0.566 0.705 0.709 0.585 0.901 0.660 0.877 0.505 0.476
Aff-R 0.748 0.782 0.803 0.905 0.685 0.735 0.806 0.851 0.841 0.887 0.647 0.760 0.535 0.447 0.422
Aff-F 0.820 0.616 0.814 0.660 0.577 0.592 0.665 0.771 0.770 0.705 0.753 0.707 0.665 0.474 0.447
Global4.8 Aff-P 0.985 0.661 1.000 0.549 0.506 0.479 0.964 1.000 0.994 0.721 0.916 0.713 0.938 0.599 0.535
Aff-R 0.922 0.867 0.876 0.749 0.933 0.737 0.865 0.889 0.875 0.907 0.912 0.996 0.900 0.894 0.554
Aff-F 0.953 0.750 0.934 0.633 0.656 0.581 0.912 0.941 0.930 0.804 0.914 0.831 0.919 0.717 0.544
Global7.2 Aff-P 0.954 0.662 0.926 0.515 0.506 0.481 0.993 0.949 0.975 0.769 0.958 0.771 0.950 0.572 0.514
Aff-R 0.935 0.857 0.893 0.939 0.937 0.648 0.836 0.928 0.881 0.861 0.895 0.993 0.866 0.872 0.510
Aff-F 0.945 0.746 0.909 0.665 0.657 0.552 0.907 0.938 0.926 0.812 0.925 0.868 0.906 0.691 0.512
Seasonal4.8 Aff-P 1.000 0.533 1.000 0.950 0.945 0.912 1.000 1.000 1.000 0.993 0.870 0.654 0.859 0.534 0.509
Aff-R 0.998 0.955 0.986 0.761 0.784 0.847 0.987 0.987 0.988 0.931 0.994 0.998 0.993 0.893 0.944
Aff-F 0.999 0.684 0.993 0.845 0.857 0.879 0.993 0.994 0.994 0.961 0.928 0.790 0.921 0.668 0.661
Seasonal7.7 Aff-P 0.992 0.534 1.000 0.628 0.970 0.882 0.998 0.985 1.000 0.982 0.868 0.697 0.925 0.482 0.546
Aff-R 0.998 0.934 0.983 0.809 0.570 0.800 0.982 0.983 0.986 0.901 0.991 0.994 0.988 0.817 0.919
Aff-F 0.995 0.679 0.992 0.707 0.718 0.839 0.990 0.984 0.993 0.940 0.925 0.820 0.956 0.606 0.685
Shapelet4.9 Aff-P 0.997 0.517 0.952 0.653 0.779 0.763 0.968 0.874 0.979 0.646 0.732 0.606 0.832 0.547 0.492
Aff-R 0.977 0.976 0.979 0.732 0.704 0.743 0.871 0.983 0.979 0.877 0.992 0.988 0.965 0.898 0.940
Aff-F 0.987 0.676 0.966 0.690 0.740 0.753 0.917 0.925 0.979 0.744 0.842 0.751 0.894 0.680 0.646
Shapelet7.4 Aff-P 0.999 0.514 0.941 0.591 0.658 0.641 0.965 0.913 0.914 0.656 0.827 0.662 0.821 0.567 0.499
Aff-R 0.969 0.975 0.972 0.842 0.659 0.823 0.967 0.970 0.974 0.943 0.985 0.983 0.949 0.870 0.873
Aff-F 0.983 0.673 0.957 0.695 0.658 0.721 0.966 0.940 0.943 0.774 0.899 0.791 0.880 0.686 0.635
Mixture5.7 Aff-P 0.863 0.628 0.867 0.555 0.541 0.802 0.875 0.867 0.581 0.779 0.560 0.539 0.879 0.543 0.501
Aff-R 0.991 0.940 0.920 0.910 0.721 0.737 0.920 0.922 0.959 0.895 0.992 0.998 0.970 0.997 0.969
Aff-F 0.923 0.753 0.893 0.689 0.618 0.768 0.897 0.894 0.723 0.833 0.716 0.700 0.922 0.703 0.660
Mixture5.9 Aff-P 0.841 0.607 0.838 0.510 0.502 0.814 0.820 0.797 0.568 0.744 0.520 0.517 0.868 0.518 0.515
Aff-R 0.988 0.871 0.991 0.954 0.608 0.718 0.989 0.990 0.975 0.929 0.991 1.000 0.947 1.000 0.995
Aff-F 0.909 0.716 0.908 0.665 0.550 0.763 0.897 0.883 0.718 0.827 0.682 0.682 0.906 0.683 0.678
Trend4.8 Aff-P 0.861 0.812 0.831 0.505 0.560 0.751 0.861 0.840 0.571 0.765 0.502 0.502 0.877 0.505 0.503
Aff-R 0.997 0.974 0.984 0.937 0.746 0.803 0.929 0.978 0.991 0.994 0.995 1.000 0.942 1.000 0.996
Aff-F 0.924 0.886 0.901 0.656 0.640 0.776 0.893 0.904 0.724 0.865 0.667 0.668 0.909 0.671 0.668
Trend7.8 Aff-P 0.809 0.716 0.707 0.522 0.609 0.656 0.722 0.770 0.582 0.712 0.516 0.519 0.865 0.517 0.506
Aff-R 0.916 0.978 0.985 0.848 0.679 0.627 0.981 0.961 0.975 0.946 0.991 1.000 0.842 1.000 0.971
Aff-F 0.859 0.827 0.823 0.647 0.642 0.641 0.832 0.855 0.729 0.812 0.679 0.684 0.853 0.681 0.665
$1^{st}$ Count 31 0 4 1 0 1 1 0 0 2 4 7 7 3 2

### v3

@@ Appendix C Full Experimental Results
Dataset Metric CATCH Modern iTrans DualTF ATrans DC TsNet Patch DLin NLin AE TFAD Ocsvm IF PCA HBOS
CICIDS Aff-P 0.667 0.563 0.586 0.553 0.531 0.533 0.566 0.569 0.575 0.579 0.543 0.550 0.530 0.548 0.541 0.538
Aff-R 0.959 0.781 0.892 0.926 0.593 0.882 0.783 0.785 0.799 0.794 0.156 0.611 1.000 0.672 0.724 0.545
Aff-F 0.787 0.654 0.708 0.692 0.560 0.664 0.657 0.660 0.669 0.669 0.243 0.579 0.693 0.604 0.619 0.542
CalIt2 Aff-P 0.742 0.650 0.703 0.617 0.645 0.571 0.691 0.667 0.668 0.616 0.560 0.619 0.652 0.539 0.688 0.620
Aff-R 0.955 0.975 0.963 0.959 0.838 0.894 0.932 0.976 0.976 0.984 0.617 0.933 0.982 0.321 0.869 0.969
Aff-F 0.835 0.780 0.812 0.751 0.729 0.697 0.794 0.793 0.793 0.757 0.587 0.744 0.783 0.402 0.768 0.756
Credit Aff-P 0.618 0.611 0.559 0.513 0.521 0.488 0.610 0.612 0.605 0.608 0.560 0.585 0.556 0.658 0.564 0.533
Aff-R 0.956 0.952 0.984 0.935 0.865 0.898 0.953 0.955 0.948 0.952 0.562 0.757 0.997 0.612 0.959 0.999
Aff-F 0.750 0.744 0.713 0.663 0.650 0.632 0.744 0.746 0.738 0.742 0.561 0.660 0.714 0.634 0.710 0.695
GECCO Aff-P 0.832 0.808 0.735 0.633 0.690 0.567 0.810 0.831 0.808 0.793 0.836 0.526 0.499 0.647 0.646 0.620
Aff-R 0.998 0.998 0.979 0.786 0.903 0.872 0.997 0.995 0.997 0.992 0.810 0.776 1.000 0.315 1.000 0.827
Aff-F 0.908 0.893 0.839 0.701 0.782 0.687 0.894 0.906 0.893 0.882 0.823 0.627 0.666 0.424 0.785 0.708
Genesis Aff-P 0.835 0.728 0.822 0.683 0.749 0.659 0.780 0.763 0.772 0.724 0.759 0.437 0.512 0.673 0.691 0.564
Aff-R 0.966 0.974 0.972 0.996 1.000 0.943 0.968 0.974 0.959 0.971 0.976 0.687 1.000 0.951 0.991 1.000
Aff-F 0.896 0.833 0.891 0.810 0.856 0.776 0.864 0.856 0.856 0.829 0.854 0.535 0.677 0.788 0.814 0.721
MSL Aff-P 0.599 0.578 0.566 0.562 0.549 0.576 0.589 0.584 0.577 0.584 0.521 0.516 0.497 0.502 0.538 0.520
Aff-R 0.966 0.975 0.951 0.618 0.933 0.874 0.973 0.952 0.975 0.948 0.781 0.936 0.902 0.697 0.914 0.982
Aff-F 0.740 0.726 0.710 0.588 0.692 0.694 0.734 0.724 0.725 0.723 0.625 0.665 0.641 0.584 0.678 0.680
NYC Aff-P 1.000 0.639 0.520 0.551 0.751 0.769 0.750 0.719 0.814 0.817 0.529 0.689 0.500 0.481 0.516 0.525
Aff-R 0.989 0.965 0.999 0.988 0.988 0.980 0.843 0.843 0.843 0.821 0.989 0.689 1.000 0.989 0.997 0.946
Aff-F 0.994 0.769 0.684 0.708 0.853 0.862 0.794 0.776 0.828 0.819 0.689 0.689 0.667 0.648 0.680 0.675
PSM Aff-P 0.808 0.734 0.765 0.622 0.600 0.538 0.762 0.739 0.777 0.762 0.776 0.543 0.652 0.904 0.712 0.621
Aff-R 0.918 0.941 0.966 0.868 0.871 0.932 0.939 0.948 0.893 0.942 0.649 0.744 0.447 0.472 0.692 0.700
Aff-F 0.859 0.825 0.854 0.725 0.710 0.682 0.842 0.831 0.831 0.843 0.707 0.628 0.531 0.620 0.702 0.658
SMD Aff-P 0.773 0.755 0.736 0.527 0.607 0.510 0.745 0.748 0.761 0.762 0.889 0.517 0.649 0.801 0.680 0.557
Aff-R 0.938 0.948 0.943 0.956 0.895 0.998 0.938 0.970 0.940 0.946 0.291 0.913 0.866 0.513 0.807 0.722
Aff-F 0.847 0.840 0.827 0.679 0.724 0.675 0.831 0.845 0.841 0.844 0.439 0.660 0.742 0.626 0.738 0.629
ASD Aff-P 0.693 0.669 0.660 0.551 0.553 0.569 0.693 0.662 0.690 0.650 0.600 0.532 0.505 0.761 0.532 0.538
Aff-R 0.966 0.953 0.969 0.760 0.870 0.921 0.963 0.954 0.918 0.944 0.942 0.785 0.810 0.842 0.871 0.891
Aff-F 0.804 0.782 0.780 0.604 0.674 0.702 0.800 0.777 0.782 0.766 0.731 0.630 0.617 0.781 0.656 0.669
Contextual4.9 Aff-P 0.838 0.505 0.821 0.503 0.517 0.490 0.552 0.669 0.773 0.563 0.861 0.526 0.619 0.861 0.488 0.487
Aff-R 0.812 0.814 0.761 0.776 0.792 0.779 0.837 0.883 0.806 0.908 0.674 0.666 0.769 0.582 0.463 0.547
Aff-F 0.825 0.623 0.790 0.611 0.625 0.602 0.666 0.761 0.789 0.695 0.756 0.588 0.686 0.694 0.475 0.515
Contextual7.2 Aff-P 0.907 0.508 0.826 0.519 0.498 0.496 0.566 0.705 0.709 0.585 0.901 0.523 0.660 0.877 0.505 0.476
Aff-R 0.748 0.782 0.803 0.905 0.685 0.735 0.806 0.851 0.841 0.887 0.647 0.580 0.760 0.535 0.447 0.422
Aff-F 0.820 0.616 0.814 0.660 0.577 0.592 0.665 0.771 0.770 0.705 0.753 0.550 0.707 0.665 0.474 0.447
Global4.8 Aff-P 0.985 0.661 1.000 0.549 0.506 0.479 0.964 1.000 0.994 0.721 0.916 0.520 0.713 0.938 0.599 0.535
Aff-R 0.922 0.867 0.876 0.749 0.933 0.737 0.865 0.889 0.875 0.907 0.912 0.679 0.996 0.900 0.894 0.554
Aff-F 0.953 0.750 0.934 0.633 0.656 0.581 0.912 0.941 0.930 0.804 0.914 0.589 0.831 0.919 0.717 0.544
Global7.2 Aff-P 0.954 0.662 0.926 0.515 0.506 0.481 0.993 0.949 0.975 0.769 0.958 0.513 0.771 0.950 0.572 0.514
Aff-R 0.935 0.857 0.893 0.939 0.937 0.648 0.836 0.928 0.881 0.861 0.895 0.577 0.993 0.866 0.872 0.510
Aff-F 0.945 0.746 0.909 0.665 0.657 0.552 0.907 0.938 0.926 0.812 0.925 0.543 0.868 0.906 0.691 0.512
Seasonal4.8 Aff-P 1.000 0.533 1.000 0.950 0.945 0.912 1.000 1.000 1.000 0.993 0.870 0.573 0.654 0.859 0.534 0.509
Aff-R 0.998 0.955 0.986 0.761 0.784 0.847 0.987 0.987 0.988 0.931 0.994 0.485 0.998 0.993 0.893 0.944
Aff-F 0.999 0.684 0.993 0.845 0.857 0.879 0.993 0.994 0.994 0.961 0.928 0.525 0.790 0.921 0.668 0.661
Seasonal7.7 Aff-P 0.992 0.534 1.000 0.628 0.970 0.882 0.998 0.985 1.000 0.982 0.868 0.788 0.697 0.925 0.482 0.546
Aff-R 0.998 0.934 0.983 0.809 0.570 0.800 0.982 0.983 0.986 0.901 0.991 0.916 0.994 0.988 0.817 0.919
Aff-F 0.995 0.679 0.992 0.707 0.718 0.839 0.990 0.984 0.993 0.940 0.925 0.847 0.820 0.956 0.606 0.685
Shapelet4.9 Aff-P 0.997 0.517 0.952 0.653 0.779 0.763 0.968 0.874 0.979 0.646 0.732 0.603 0.606 0.832 0.547 0.492
Aff-R 0.977 0.976 0.979 0.732 0.704 0.743 0.871 0.983 0.979 0.877 0.992 0.512 0.988 0.965 0.898 0.940
Aff-F 0.987 0.676 0.966 0.690 0.740 0.753 0.917 0.925 0.979 0.744 0.842 0.554 0.751 0.894 0.680 0.646
Shapelet7.4 Aff-P 0.999 0.514 0.941 0.591 0.658 0.641 0.965 0.913 0.914 0.656 0.827 0.738 0.662 0.821 0.567 0.499
Aff-R 0.969 0.975 0.972 0.842 0.659 0.823 0.967 0.970 0.974 0.943 0.985 0.907 0.983 0.949 0.870 0.873
Aff-F 0.983 0.673 0.957 0.695 0.658 0.721 0.966 0.940 0.943 0.774 0.899 0.814 0.791 0.880 0.686 0.635
Mixture5.7 Aff-P 0.863 0.628 0.867 0.555 0.541 0.802 0.875 0.867 0.581 0.779 0.560 0.620 0.539 0.879 0.543 0.501
Aff-R 0.991 0.940 0.920 0.910 0.721 0.737 0.920 0.922 0.959 0.895 0.992 0.884 0.998 0.970 0.997 0.969
Aff-F 0.923 0.753 0.893 0.689 0.618 0.768 0.897 0.894 0.723 0.833 0.716 0.729 0.700 0.922 0.703 0.660
Mixture8.9 Aff-P 0.841 0.607 0.838 0.510 0.502 0.814 0.820 0.797 0.568 0.744 0.520 0.561 0.517 0.868 0.518 0.515
Aff-R 0.988 0.871 0.991 0.954 0.608 0.718 0.989 0.990 0.975 0.929 0.991 0.897 1.000 0.947 1.000 0.995
Aff-F 0.909 0.716 0.908 0.665 0.550 0.763 0.897 0.883 0.718 0.827 0.682 0.690 0.682 0.906 0.683 0.678
Trend4.8 Aff-P 0.861 0.812 0.831 0.505 0.560 0.751 0.861 0.840 0.571 0.765 0.502 0.571 0.502 0.877 0.505 0.503
Aff-R 0.997 0.974 0.984 0.937 0.746 0.803 0.929 0.978 0.991 0.994 0.995 0.553 1.000 0.942 1.000 0.996
Aff-F 0.924 0.886 0.901 0.656 0.640 0.776 0.893 0.904 0.724 0.865 0.667 0.562 0.668 0.909 0.671 0.668
Trend7.8 Aff-P 0.809 0.716 0.707 0.522 0.609 0.656 0.722 0.770 0.582 0.712 0.516 0.628 0.519 0.865 0.517 0.506
Aff-R 0.916 0.978 0.985 0.848 0.679 0.627 0.981 0.961 0.975 0.946 0.991 0.848 1.000 0.842 1.000 0.971
Aff-F 0.859 0.827 0.823 0.647 0.642 0.641 0.832 0.855 0.729 0.812 0.679 0.721 0.684 0.853 0.681 0.665
$1^{st}$ Count 35 0 4 1 1 1 1 0 0 2 5 0 10 7 0 2
Appendix D Empirical verification shows that CATCH can perform fine-grained modeling in each frequency band
We design a set of experiments to validate the advantages of our method in capturing fine-grained frequency characteristics. We use the synthetic dataset creation method provided by TODS (Lai et al., 2021) to generate four synthetic datasets—see Figure 7, simulating scenarios with known frequency-specific anomalies: low-frequency anomalies, medium-frequency anomalies, high-frequency anomalies, and anomalies that are not distinctly separated across different frequency bands. We then test CATCH and several state-of-the-art methods on these datasets.
The experimental results in Table 11 show that CATCH consistently outperforms other algorithms across all four datasets. Notably, most algorithms perform poorly when dealing with time series containing high-frequency anomalies, while CATCH demonstrates outstanding performance. Moreover, even on time series where anomalies are uniformly distributed across each frequency band, CATCH still achieves outstanding results. These findings provide strong evidence that CATCH excels in capturing fine-grained frequency characteristics, significantly enhancing the performance of time series anomaly detection.
(a) Time series containing low-frequency anomalies
(b) Time series containing medium-frequency anomalies
(c) Time series containing high-frequency anomalies
(d) Anomalies that are not distinctly separated across different frequency bands
Figure 7: Four synthetic time series simulating scenarios with known frequency-specific anomalies.
Table 11: Average A-R (AUC-ROC) and Aff-F (Affiliated-F1) accuracy measures for 4 synthetic datasets of different types of anomalies. The best results are highlighted in bold, and the second-best results are underlined.
Model CATCH Modern iTrans DualTF TFAD
Metric Aff-F A-R Aff-F A-R Aff-F A-R Aff-F A-R Aff-F A-R
Low-frequency anomalies 0.998 0.996 0.811 0.873 0.850 0.929 0.744 0.662 0.700 0.538
Mid-frequency anomalies 0.927 0.987 0.814 0.859 0.913 0.984 0.723 0.575 0.638 0.519
High-frequency anomalies 0.915 0.987 0.738 0.809 0.864 0.927 0.462 0.602 0.668 0.513
Not distinctly frequency band 0.873 0.917 0.775 0.765 0.852 0.910 0.720 0.630 0.619 0.495